how do i figure out the sequence of 7, 3, -1, -5
thank you amy

Answer by Gogonati(855)

We see the difference between two successive terms:
3-7=-4, -1-3=-4, -5-(-1)=-4. The difference of this sequence is -4
the sequence will be:7, 3, -1, -5, -9, -13, -17, ...